A CAR TRAVELS THE FIRST THIRD OF TOTAL DISTANCE AT A SPEED OF V1 m/s .  THE REMAINING DISTANCE IS COVERED SUCH THAT FIRST HALF TIME IT TRAVELS WITH V2 m/s AND SECOND HALF TIME WITH V3 m/s. FIND THE AVERAGE SPEED. PLZ EXPLAIN.
 

Grade:9

2 Answers

Ayush Anand
5 years ago
let total distance be x.
distance travelled with v1 velocity=x/3
distance travelled with v2 velocity =\frac{\frac{2x}{3} }{2}
=\frac{2x}{6}
=\frac{x}{3}
distance travelled with v2 velocity =same distance as of v2 velocity
=\frac{x}{3}
So, average speed=\frac{v1+v2+v3}{3}
 
=\frac{\frac{x}{3}+\frac{x}{3}+\frac{x}{3}}{3}
=\frac{\frac{3x}{3}}{3}
=\frac{x}{3}
Ashish
5 years ago
Let the total distance=x
Distance travelled with velocity V1=x/3
".                   ".              ".       ".      V2=x/6
".                   ".                  ".        V3=.  x/12
Total time = 7x/12 and total distance= x
Now, average speed= 12/7 ms^-1